OpenStudy (anonymous):

Approximately 15% of the calls to an airline reservation phone line result in a reservation being made. (a) Suppose that an operator handles 10 calls. What is the probability that none of the 10 calls result in a reservation? (Give the answer to 3 decimals places.)

OpenStudy (anonymous):

probability of a call resulting in a reservation = .15 probability of a call not resulting in a reservation = .85 15 calls are made. what is the probability that at least 1 of them will be a reservation? that equals 1 minus the probability that all of them do not result in a reservation. since the probability of any one call not resulting in a reservation equals .85, then the probability that all 15 calls will not result in a reservation equals .85^15. that number is .087354219. the probability that at least 1 call will result in a reservation is 1 minus the probability that no calls will result in a reservation. that probability = .912645781. that's the probability that at least 1 call will result in a reservation.

OpenStudy (anonymous):

but it's 10 calls?

OpenStudy (anonymous):

(a) Suppose that an operator handles ***10 calls***. What is the probability that none of the 10 calls result in a reservation? (Give the answer to 3 decimals places.)

OpenStudy (anonymous):

then take my equation and do it for 10 calls?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

What's part C, then ' (c) What is the probability that at least one call results in a reservation being made? (Give the answer to 3 decimals places.)'

OpenStudy (aum):

\[ \large P(\text{ k successes in n trials }) = \left(\begin{matrix}n \\ k\end{matrix}\right)p^{k}q^{n-k} \\ \large n = 10;~~ k = 0; ~~p = 0.15; ~~q = 0.85 \]

OpenStudy (anonymous):

what does that mean?

OpenStudy (anonymous):

what do I have to do with the n and k? Multiply them? Divide?

OpenStudy (anonymous):

Within the parenthesis ^

OpenStudy (aum):

Have you done binomial probability distribution yet? \[ \left(\begin{matrix}n \\ k\end{matrix}\right) = \frac{n!}{(n-k)! * k!} \]

OpenStudy (anonymous):

thank you! And no...